We had to move our flipper mirror to share the beam between Peter's setup and ours as our flipper is at the place where the ISS PD array base is supposed to be!
There was no place to insert the flipper in the setup. We (Peter and Koji) decided to move the laser back for ~2".
This entirely changed the alignment of the setup. The fiber coupler was my reference of the alignment.
Once the beam is aligned, I check the coupling to the fiber. It was 50%.
I tweaked the lens and eventually the coupling is improved to 83%. (24.7mW incident, 20.4mW obtained.)
Then, I started to check the AOM path. I noticed that the 1st (or -1st) order beam is very weak.
The deflection efficiency is ~0.1%. Something is wrong.
I checked the driver. The driver's coupler output (1:10) show the amplitude ~1V. (good)
I check the main output by reducing the offset. When the coupler output is 100mV, the main output was 1V. (good)
So is the AOM itself broken???